# LLM-Generated Black-box Explanations Can Be Adversarially Helpful

### v2

@@ 7 A structural analysis towards adversarial helpfulness
Here we dive deeper into the mechanism of the adversarial helpfulness phenomena that we observe in previous sections. As Figure 3 illustrates, the explanation in natural language has an inherent structure that resembles a graph.
The literature about discourse analysis and automatic reasoning has drawn analogies between reasoning, explanation (and discourse in general) and graphs. One of the seminal works along this direction is Rhetorical Structure Theory (Mann & Thompson, 1988), which identified spans of texts (discourse units) as graph nodes and specified the discourse relations as graph edges. ConceptNet (Liu & Singh, 2004) and other knowledge graphs specified concepts as graph nodes, and abstracted the relations as graph edges. The ECQA dataset (Aggarwal et al., 2021b) that we use is based on Commonsense QA (Talmor et al., 2019), which is based on ConceptNet. Dziri et al. (2024) abstracted the compositional reasoning problems into graphs while studying the difficulty of the reasoning problems. Prystawski et al. (2024) used a Bayesian network to model how the reasoning emerges from the locality of experience. CLEAR (Ma et al., 2022) and RSGG-CE (Prado-Romero et al., 2024) leveraged graph structures to generate counterfactual explanations. Following their avenues of research, we set up a graph-based symbolic reasoning problem as an abstraction of the “explanation towards the incorrect answer” phenomenon.
Figure 3: Explanation has a graph-like structure, where each graph node is analogous to a “discourse unit” and each graph edge serves as a “reasoning step”.
@@
We consider the process of explanation to be an instance of path finding on a graph. In this path-finding problem, we find a path from the “root” node to the “answer” node. The explanation serves as the path that connects the known (the “root” node) to the unknown (the “answer” node).
Figure 4: An example of symbolic reasoning graph. Left: the graph appearing in DOT string format. Right: the graph appearing in graph format. If the path “root $\rightarrow$ 0_1 $\rightarrow$ 0_2 $\rightarrow$ A” is the reasoning path supporting answer A, supporting answer C would need a reasoning path “root $\rightarrow$ 2_1 $\rightarrow$ 2_2 $\rightarrow$ C”.
Following specifies how the graphs are constructed. There are two parameters: number of branches $B$ and path length $L$.
@@
2. We specify $B$ branches. All branches start from the root node, and extends by $L$ steps. The $j^{th}$ node at the $i^{th}$ branch is marked $i_{j}$, and the last node at each branch is marked with an alphabet (A, B, C, and so on), to resemble the answer in the multiple-choice question.
LLMs have limited capabilities in capturing the structural information in graphs (Huang et al., 2023a). Based on the intuition, we attach two simplification assumptions. First, all branches have the same path length. Second, each non-root node uniquely appears on only one path. Figure 4 shows an example of such graph with $B=3$ branches of $L=4$ path.
3. We linearize each graph using the DOT format. Together with this graph string and a brief description of the formatting specifications, we prompt the LLM to find the alternative answer, and return us with the path from root to the alternative answer. The correctness of the returned path is evaluated with exact match. [^4] Empirically we find that many LLMs tend to write the last node twice. For example, for the graph in Figure 3, instead of the reasoning path “root $\rightarrow$ 2_1 $\rightarrow$ 2_2 $\rightarrow$ C”, the LLMs sometimes write “root $\rightarrow$ 2_1 $\rightarrow$ 2_2 $\rightarrow$ 2_3 $\rightarrow$ C”. We consider this correct as well, if we allow the $\rightarrow$” to signal both a graph reasoning step and a name aliasing step.
4. We compute the success rate of alternate path finding at a given graph complexity. To reduce the complexity of experiments, we fix $B=L$, and take this number as the “graph complexity”. At a given graph complexity $B$, there are $B\times(B-1)$ alternate path-finding cases. Among them, the portion of correctly returned path is defined as the success rate.
@@
An LLM might generate responsed based on the naming patterns instead of the graph path structure. To deconfound this factor, we run another version of symbolic graph reasoning experiment. This time, we replace the node names of each non-leaf and non-root node with a randomly-chosen but non-overlapping character.
@@ Results
Figure 5 plots the success rate against the number of graph branches, without (left) and with (right) the randomization step of the node names. As the complexity of the graph increase, the success rate of alternative path finding decreases. When we factor out the reliance on the node names, the performances of all models (except GPT-4/4 Turbo) drop to zero. We consider one graph node to be the abstraction of one sentence, so the complexity of the graphs resemble that of the generated explanations. Why, then, do the LLM-generated explanations show adversarial helpfulness while these LLMs fail badly at handling the structures of the adversarial helpfulness? Apparently, two factors jointly contribute to the “adversarial helpfulness”: the structure and the lexical contents. When an LLM is not capable of handling the structures, it can still utilize the lexical contents to produce adversarially helpful explanations.
@@
8 Discussion
Adversarial helpfulness vs unfaithfulness
There are intricate differences between these two improvement areas of LLM-generated explanations. Adversarial helpfulness refers to the explainer’s behavior that rationalizes an incorrect problem. (Un)faithfulness, on the other hand, is not tied to the correctness of the problem to be explained. For example, an explanation for a decision-making model can be unfaithful if the stated reasons differ from the reasoning process of this model.
Adversarial helpfulness vs plausibility
Plausibility is different from faithfulness — for an extended discussion of their difference please refer to Agarwal et al. (2024). Here we consider that adversarial helpfulness overlaps with plausibility; the latter indicates a potential negative impact on humans, whereas the former does not. Plausibility is a feature, but adversarial helpfulness is a bug.
Adversarial helpfulness vs hallucination
Hallucination refers to the undesirable phenomena of natural language generation (NLG) systems generating unfaithful or nonsensical texts (Ji et al., 2023). Adversarial helpfulness describes phenomena from a much smaller scope: those where the explanations facilitate the belief of the incorrect explanandum.
Do the guardrails work?
When trying to let LLMs produce adversarially helpful explanations on the reported datasets, we find that the existing guardrails are very weak, if they exist at all. Many existing guardrails rely on the models’ parametric knowledge (“common-sense”) to avoid outputting adversarial explanations. This mechanism can be rewritten by simply asking the LLM to, e.g., “hypothesize an alternative scenario”. Since we should not prevent an LLM from reasoning counterfactually, one can argue that the adversarial helpfulness problem is unable to be guardrailed at the LLM level. Instead, this problem should be guarded at the user level.
Safe use of AI as explainers
We provide three recommendations here.
First, let humans make decisions, not AIs. We can use the AI explainer as a “Prudence” model (Miller, 2023). A “Prudence” model provides evidence supporting human’s decisions, without directly giving us an answer. The alternative, “Bluster” model, recommends an answer, optionally with the rationales supporting that answer. Since the rationales supporting the answer can be adversarially helpful, it is important to delegate the least possible amount of “decision-making” responsibility to the explainers (especially those based on generative AI). [^5] Also note that we should avoid overloading the human users with information about the model, since exposing too much model-related details could make humans less able to detect models’ mistakes (Poursabzi-Sangdeh et al., 2021).
Second, instruct AIs to generate rationales that support multiple alternative answers to offset the cherry-picking behavior of AI explainers.
Third, pass in as much of the decision-maker model’s intermediate signals as possible, especially when the decisions are difficult (i.e., the decision-maker model can likely produce an incorrect label). Note that LLMs still struggle at summarizing the neurons’ activations (Huang et al., 2023b) — perhaps because the neurons are too fine-grained (Niu et al., 2024) — but this struggle should not prevent us from giving the explainer a guideline to follow.
9 Conclusion
We identify a potentially perilous scenario of using LLMs as explaining assistants in a “black-box” manner, which we call adversarial helpfulness. When prompted to explain an incorrect answer, LLMs can generate convincing explanations, making wrong answers look right. We show that this issue affects both humans and LLM evaluators. We analyze the persuasion strategies, and find that LLMs frequently reframe the questions and present selective details, among other strategies, in favor of the incorrect answers. We further analyze the LLMs’ mechanisms by setting up a symbolic reasoning problem. The findings motivate us to recommend future practices for using LLMs as explainers.
Appendix A Annotation User Interface
@@
Appendix B GPT-4 Turbo prompt for identifying the persuasion strategies
@@
Appendix C Additional results for persuasion strategy identification

### v3

@@ 10. Complex inference 4 1 4 9 8 5 3 7 3
6 Strategies toward adversarial helpfulness
Recent literature involves many taxonomies of persuasion strategies. For example, Dimitrov et al. (2021) identified strategies in social media texts, Piskorski et al. (2023) considered news, and Zeng et al. (2024) considered 40 techniques for jailbreaking as well as other NLP tasks. Inspired by these works, we identify ten persuasion strategies that are particularly relevant to LLM-generated explanations. A brief summary of each strategy is included in Appendix C.
We use one of the LLMs with the strongest reasoning abilities, GPT-4 Turbo (gpt-4-0125-preview), to detect the persuasion strategies used in the explanations. Appendix C lists the prompt. If no strategy is detected, the json object parsed from the LLM’s response would contain zero in all ten entries. Figure 1 lists two examples of the identified strategies.
Frequencies of the strategies
First, the sheer frequencies themselves are alarming. For the commonsense questions, more than $70\%$ of the explanations highlight the comparative advantage towards incorrect answers. For inference tasks, more than $90\%$ of the explanations involve reframing the questions. While these results are only for the commonsense and inference datasets, LLMs are capable of including these persuasion strategies when explaining questions in real-world scenarios. We expect that LLMs also exhibit these strategies in correct cases, since these persuasion capabilities would still exist. However, we argue that a safe explainer should minimize the use of these persuasion strategies in explanation, especially when the explanandum involves an incorrect problem.
Second, the three explainer models show common trends in applying persuasion strategies. The LLM-generated explanations demonstrate elevated confidence levels. [^2] This may related to the overly-confident tone in our prompts, as listed in Figure 2. Strategies like selective evidence, and selective fact presentation are frequently used. [^3] Relatedly, selectivity is a desirable feature in human explanations (Lombrozo and Liquin, 2023). The strategies like ‘appeal to authority’ and ‘analogical evidence’ are infrequently used in any of the models. These indicate that adversarial helpfulness could largely be safeguarded by defending only a finite set of persuasion strategies.
For completeness, we repeat the automatic detection experiments using the taxonomy of Zeng et al. (2024) – the experiment results are shown in Table 3 of the Appendix. It shows similar observations that only a few of the persuasion strategies are applied (e.g., logical appeal, encouragement, and framing), but very frequently.
@@ 7 A structural analysis towards adversarial helpfulness
Figure 3: Two explanations towards two answer choices for an ECQA problem, where each graph node is analogous to a reasoning unit, and each graph edge serves as a reasoning step.
Here we provide a graphical inquiry into the mechanism of the adversarial helpfulness phenomena that we observe in previous sections. As Figure 3 illustrates, the explanation in natural language has an inherent graph structure.
The literature on discourse analysis and automatic reasoning has drawn analogies between reasoning, explanation (and discourse in general), and graphs. One of the seminal works in this direction is Rhetorical Structure Theory (Mann and Thompson, 1988), which identified spans of texts (discourse units) as graph nodes and specified the discourse relations as graph edges. ConceptNet (Liu and Singh, 2004) and other knowledge graphs specified concepts as graph nodes, and abstracted the relations as graph edges. The ECQA dataset (Aggarwal et al., 2021b) that we use is based on Commonsense QA (Talmor et al., 2019), which is based on ConceptNet. Dziri et al. (2024) abstracted the compositional reasoning problems into graphs while studying the difficulty of the reasoning problems. Prystawski et al. (2024) used a Bayesian network to model how reasoning emerges from the locality of experience. CLEAR (Ma et al., 2022) and RSGG-CE (Prado-Romero et al., 2024) leveraged graph structures to generate counterfactual explanations. Following these avenues of research, we set up a graph-based symbolic reasoning problem as an abstraction of the “explanation towards the incorrect answer” phenomenon.
@@
We consider the process of explanation to be an instance of path finding on a graph. In each problem, we find a path from the root node to a leaf node. The explanation serves as the path that connects the problem (the root node) to the answer (the leaf node).
Figure 4: Left: Example of a symbolic reasoning graph with non-randomized node names. Right: Example of a symbolic reasoning graph with randomized node names. The graph in string format, the graph plotted. If the path “root $\rightarrow$ 0_1 $\rightarrow$ 0_2 $\rightarrow$ A” is the reasoning path supporting answer A, supporting answer C would need a reasoning path “root $\rightarrow$ 2_1 $\rightarrow$ 2_2 $\rightarrow$ C”.
The following specifies how the graphs are constructed. There are two parameters: number of branches $B$ and path length $L$.
@@
2. We specify $B$ branches. All branches start from the root node, and extend by $L$ steps. The $j^{th}$ node at the $i^{th}$ branch is marked $i_{j}$, and the last node at each branch is marked with an alphabet (A, B, C, and so on), to resemble the answer in the multiple-choice question.
LLMs have limited capabilities in capturing the structural information in graphs (Huang et al., 2023a). Based on the intuition, we attach two simplification assumptions. First, all branches have the same path length. Second, each non-root node uniquely appears on only one path (effectively making it a tree). Figure 4 shows an example of such a graph with $B=3$ branches of $L=4$ path.
3. We linearize each graph. Together with this graph string and a brief description of the formatting specifications, we prompt the LLM to find the alternative answer with the supporting path. The correctness of the returned path is evaluated with exact match. [^4] Empirically, we find that many LLMs tend to write the last node twice. For example, for the graph in Figure 3, instead of the reasoning path “root $\rightarrow$ 2_1 $\rightarrow$ 2_2 $\rightarrow$ C”, the LLMs sometimes write “root $\rightarrow$ 2_1 $\rightarrow$ 2_2 $\rightarrow$ 2_3 $\rightarrow$ C”. We consider this correct as well, if we allow the $\rightarrow$” to signal both a graph reasoning step and a name aliasing step. We compute the success rate of alternate path finding at a given graph complexity. To reduce the complexity of experiments, we fix $B=L$, and take this number as the “graph complexity”. At a given graph complexity $B$, there are $B\times(B-1)$ alternate path-finding cases. The success rate is the portion of the correctly returned path among them.
@@
An LLM might generate responses based on the naming patterns instead of the graph path structure. For example, one such pattern is considering the nodes “0_0” and “0_1” to be connected regardless of the real connectivity. To deconfound such bias, we run another version of the symbolic graph reasoning experiment. This time, we replace the node names of each non-leaf and non-root node with a randomly chosen but non-overlapping character.
Example
Consider a graph with non-random node names:
  • Path to answer A: root $\rightarrow$ 0_1 $\rightarrow$ 0_2 $\rightarrow$ 0_3 $\rightarrow$ A
  • Path to answer C: root $\rightarrow$ 2_1 $\rightarrow$ 2_2 $\rightarrow$ 2_3 $\rightarrow$ C
Here, the LLM might rely on patterns like "0_1" to "0_3".
Now, with randomized node names:
  • Path to answer A: root $\rightarrow$ ] $\rightarrow$ P $\rightarrow$ % $\rightarrow$ A
  • Path to answer C: root $\rightarrow$ | $\rightarrow$ 2 $\rightarrow$ = $\rightarrow$ C
The LLM must understand the actual graph connections rather than relying on familiar patterns.
@@ Results
Figure 5 of Appendix A plots the success rate against the number of graph branches, without (left) and with (right) the randomization step of the node names. As the complexity of the graph increases, the success rate of alternative path finding decreases. When we factor out the reliance on the node names, the performances of all models (except GPT-4/4 Turbo) drop to zero. Even the highest-performing model, GPT-4, fails in nearly half of the graphs with only $L=6$. LLMs, including GPT-4, struggle with more complex graphs (higher B), suggesting limitations in their reasoning abilities when pathfinding in structured data. Recall that each graph reasoning step is an abstraction of a sentence, a path with length $L=6$ represents an explanation with reasonable complexity. Therefore, one might find the high failure rates surprising. We hypothesize that multiple factors jointly contribute to the “adversarial helpfulness”, including at least the explanation structures and the lexical contents. When an LLM cannot handle the structures, it can still use the lexical contents to produce adversarially helpful explanations. The results demonstrates that LLMs, including GPT-4, tend to rely on superficial naming patterns rather than comprehending the actual graph structure, leading to misleading explanations; this is evidenced by the significant drop in success rates when node names are randomized, revealing the models’ limitations in reasoning and emphasizing the need for improved methods to ensure reliable and accurate explanations.
8 Discussion
Let us first contrast adversarial helpfulness (AH) with other terms commonly used to describe the pitfalls of explanation.
vs unfaithfulness
There are subtle differences between these two concerns. AH refers to the explainer’s behavior that rationalizes an incorrect problem. (Un)faithfulness, however, is not tied to the correctness of the explained problem and answer.
vs plausibility
AH overlaps with plausibility, but there are distinctions. An explanation is plausible if it is coherent with human reasoning and understanding (Agarwal et al., 2024). An AH explanation is not necessarily coherent, but could potentially be manipulative to human reasoning and understanding. Plausibility is a feature, but AH is a bug.
vs hallucination
Hallucination refers to the undesirable phenomena of natural language generation (NLG) systems generating unfaithful or nonsensical texts (Ji et al., 2023). AH describes phenomena in a much smaller scope: those where the explanations facilitate the belief of the incorrect explanandum.
vs sycophancy
Sycophancy refers to model responses that match user beliefs (even if they are not truthful) (Sharma et al., 2023). AH explanations do not necessarily match user beliefs. Instead of being untruthful, these explanations usually present truths selectively.
vs overtrust / over-reliance / miscalibration
AH describes a property in the explanations, whereas these describe the behavior or states of the human users.
Existing LLM guardrails cannot defend against adversarial helpfulness
When trying to let LLMs produce AH explanations on the reported datasets, we find that the existing guardrails are very weak, if they exist at all. We consider the reason to be that the existing guardrails do not inhibit the models’ abilities that facilitate many AH strategies, e.g., ignoring “unimportant” facts, stating explanations confidently, and (re-)framing the problems for easier understanding. Since these abilities are crucial for LLMs to function normally, we make an even bolder claim here: that AH cannot be fully guardrailed at the LLM level. Instead, this problem should be guarded at the user level. In other words, the developers of LLM explainers should avoid using LLMs to explain incorrect problems.
Safe use of LLM-based explainers
We provide three recommendations here.
First, delegate the least possible amount of decision-making responsibility to AI explainers. We can use the AI explainer as a “Prudence” model (Miller, 2023) which provides evidence supporting human decisions, without directly giving us an answer. The alternative, “Bluster” model (Miller, 2023), recommends an answer, optionally with the rationales supporting that answer, but the rationale can be adversarially helpful. [^5] Also note that we should avoid overloading the human users with information about the model since exposing too many model-related details could make humans less able to detect models’ mistakes (Poursabzi-Sangdeh et al., 2021).
Second, instruct AIs to generate rationales supporting multiple alternative answers to offset their selective presentation of facts, a frequently identified AH strategy.
Third, pass in as much of the decision-maker model’s intermediate signals as possible, especially when the decisions are difficult (i.e., the decision-maker model can likely produce an incorrect label). Note that LLMs still struggle at summarizing neuron activations (Huang et al., 2023b) — perhaps because the neurons are too fine-grained (Niu et al., 2024) — but this struggle should not prohibit passing the model intrinsics to the explainer.
9 Conclusion
We identify a potentially perilous scenario, which we call ‘adversarial helpfulness’, that arises from using LLMs as explanation assistants in a “black-box” manner. When prompted to explain an incorrect answer, LLMs can generate convincing explanations, making incorrect answers look correct. We show that this issue affects both humans and LLM evaluators. We analyze the persuasion strategies, and find that LLMs frequently reframe the questions and present selective details, among other strategies, in favor of the incorrect answers. We set up a symbolic graph reasoning problem as an abstract of adversarial helpful explanations, and find that the LLMs rely more on the lexical cues than the discourse structures. The findings motivate us to recommend future practices for using LLMs as explainers.
10 Limitations
Variances in item-wise results
When getting down to the item-wise level, the evaluators show varying trends. First, humans correlate weakly with proxy models. On the dataset where human annotator results are available, we compute the Pearson correlations between the averaged human results and the evaluator models. None of the correlations are significant, indicating that the evaluator models show very different fluency, correctness, and convincingness ratings from the human annotators. Second, human results show poor inter-annotator agreement. This is because the MTurk platform distributes the annotation tasks to more than three annotators. Third, the proxy models assign different scores. We compute the Cohen’s Kappa between each pair of the proxy models. None of them have a value larger than 0.1 for any score, indicating poor agreement between the evaluators. This is because the proxy models have different “baseline marking guidelines”, as is illustrated by the drastically different mean scores in Table 1.
Human evaluations
Regarding human evaluations, some additional studies could test specific biases, e.g., whether uninformative explanations can improve the convincingness ratings.
Persuasion strategies
We present an exploratory analysis of the explanation strategies, opening up future research directions. First, the cause of each strategy can be analyzed by, for example, correlating each persuasion strategy and linguistic marker, like syntactic complexity. Second, how each of the strategies affects the adversarial helpfulness can be studied in future work.
Appendix A Result plots
@@
Appendix B Annotation User Interface
@@
Appendix C GPT-4 Turbo prompt for identifying the persuasion strategies
@@
Appendix D Additional results for persuasion strategy identification